OpenStudy (anonymous):

What is the value of g[f(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? (Type only the value of your final answer, no variables.)

OpenStudy (anonymous):

morning saifoo!!!

OpenStudy (saifoo.khan):

Morning!! did u sleep well?

OpenStudy (anonymous):

i did thank u :D

OpenStudy (lalaly):

\[f(5)=10+1=11\]\[g(f(5))=g(11)=22-5=17\]

OpenStudy (saifoo.khan):

2x - 5 2(2x+1) -5 2(2(5)+1) - 5 2(11) -5 22 - 5 17